# AdaMesh: Personalized Facial Expressions and Head Poses for Adaptive Speech-Driven 3D Facial Animation

### v3

@@ Abstract
Speech-driven 3D facial animation aims at generating facial movements that are synchronized with the driving speech, which has been widely explored recently. Existing works mostly neglect the person-specific talking style in generation, including facial expression and head pose styles. Several works intend to capture the personalities by fine-tuning modules. However, limited training data leads to the lack of vividness. In this work, we propose AdaMesh, a novel adaptive speech-driven facial animation approach, which learns the personalized talking style from a reference video of about 10 seconds and generates vivid facial expressions and head poses. Specifically, we propose mixture-of-low-rank adaptation (MoLoRA) to fine-tune the expression adapter, which efficiently captures the facial expression style. For the personalized pose style, we propose a pose adapter by building a discrete pose prior and retrieving the appropriate style embedding with a semantic-aware pose style matrix without fine-tuning. Extensive experimental results show that our approach outperforms state-of-the-art methods, preserves the talking style in the reference video, and generates vivid facial animation. The supplementary video and code will be available at https://adamesh.github.io.
1 Introduction
@@
Speech-driven 3D facial animation has been widely explored and attracted increasing interest from both academics and industry. This technology has great potential in virtual reality, film production, and game creation. Most previous works focus on improving synchronization between speech and lip movements (Cudeiro et al., 2019; Richard et al., 2021; Fan et al., 2022; Xing et al., 2023), however, they have neglected the personalized talking style, including the facial expressions and head poses. The facial animation with expressive facial expressions and diverse head poses contributes to a more vivid virtual character and grants users with a better interaction experience.
@@
Recently, some works (Ye et al., 2023; Thambiraja et al., 2023) attempt to model the person-specific talking style by fine-tuning or adapting modules. In the real application, only a limited amount of video clips, even shorter than 1 minute, are provided by target users to capture the personalized talking style. These methods thus meet several challenges. 1) The few adaptation data probably induces catastrophic forgetting (Zhai et al., 2023) for the pre-trained model and easily leads to the overfitting problem (Zhang et al., 2022). Specifically for the facial expressions, after the adaptation, the lip synchronization and richness of expressions significantly decrease for the unseen speech inputs. 2) Speech is a weak control signal for the head poses (Henter et al., 2020; Zhang et al., 2021). Adaption with few data or learning mappings on such weak signals leads to averaged generation. The predicted head poses are in lack of diversity and have a smaller variation of movements.
@@
In this work, to address these two challenges, we propose an adaptive speech-driven 3D facial animation approach, AdaMesh, consisting of an expression adapter and a pose adapter. To tackle the catastrophic forgetting and overfitting problem, we devise an expression adapter built upon the low-rank adaptation technique, LoRA(Hu et al., 2022), for its high data efficiency in few-shot model adaptation. Since the facial expressions contain temporal dynamics, we propose a novel mixture-of-LoRA (MoLoRA) strategy, featured by employing LoRA with different rank sizes to model the multi-scale dynamic facial details.
@@
As for the averaged pose generation problem, we recognize it as a repercussion of generating poses via motion regression (Xing et al., 2023) and a mismatch of the adapted pose style. Hence we formulate the pose generation task as a generative one, for which we employ a vector quantized-variational autoencoder (VQ-VAE) (van den Oord et al., 2018) and a generative pre-trained Transformer (Pose GPT) network (Vaswani et al., 2017) for the backbone of the pose adapter. The head poses are mostly associated with the conveyance of negation, affirmation, and turnaround in semantics. To consider and balance both strong and weak semantic associations in head pose patterns, we propose the retrieval strategy for the pose style adaptation with a novel semantic-aware pose style matrix.
The main contributions of our work can be summarized as follows:
@@
  • We propose AdaMesh, which is the first work that takes both facial expressions and head poses into account in personalized speech-driven 3D facial animation.
@@
  • We are the first to introduce LoRA to the 3D facial animation task and propose MoLoRA to efficiently capture the multi-scale temporal dynamics in facial expressions from limited data.
@@
  • We propose a pose adapter, featured by a semantic-aware pose style matrix and a simple retrieval strategy for more diverse pose generation without fine-tuning any parameters.
@@
2 Related Work
@@
Speech-Driven 3D facial Animation. The recent speech-driven 3D facial animation methods can be divided into two categories according to the format of animation. The first category of methods (Cudeiro et al., 2019; Richard et al., 2021; Fan et al., 2022; Xing et al., 2023; Thambiraja et al., 2023) directly predicts vertices of the face from speech and renders the vertices to meshes for visualization. The vertice data is collected through professional 3D scans of human heads (Fanelli et al., 2010; Cudeiro et al., 2019; Li et al., 2017), which is expensive and time-consuming. The scanned data precisely describes the facial movements. The methods trained on such data usually have authentic lip movements. Another category of methods predicts parameters of expressions (Daněček et al., 2023) or blendshapes (Chen et al., 2022; Peng et al., 2023). These parameters can be reconstructed from the scanned vertices or 2D face videos through a parametric face model (Li et al., 2017; Peng et al., 2023). However, the decreasing quality caused by the reconstruction usually leads to lower speech-lip synchronization. The latest methods that take personalized style into consideration mostly adopt expression parameters for style modeling, since the scanned vertices are in the neutral style and lack rich expressions. Emotalk (Peng et al., 2023) and EMOTE (Daněček et al., 2023) are related works, which disentangle content and emotion to generate emotional talking style. But they adopt private or unavailable reconstruction pipelines. The most related work is Imitator (Thambiraja et al., 2023), which generates meshes and leverages the expressions from 2D video for style adaptation.
@@
Head Poses Generation. To the best of our knowledge, existing speech-driven 3D facial animation methods neglect the head poses, since the scanned vertices have no head movements. Some portrait-realistic talking face generation methods attempt to model the head poses. A typical way is to copy the head pose sequence from a real video recording (Thies et al., 2020; Ye et al., 2023; Chen et al., 2023), which can definitely generate realistic head poses. However, these methods disregard the correlation between speech and head pose, and the repetitive and non-engaging head poses create unnaturalness during long periods of interaction. The rapidly growing research topic, gesture generation from speech (Zhu et al., 2023), can also justify the importance of predicting head poses from speech rather than utilizing a template pose sequence. Another way is to predict head poses from speech (Lu et al., 2021; Zhang et al., 2021). The most effective method for head poses is FACIAL (Zhang et al., 2021), which proposes an implicit attribute learning framework to generate natural head poses. However, FACIAL sometimes generates static head poses and lacks pose diversity. Moreover, these methods can only generate head poses for a specific person, and cannot realize flexible style adaptation. Our approach could be regarded as bridging the advantages of the aforementioned two ways. The quantized pose space learned by VQ-VAE ensures realism, while the prediction step outputs natural but not repetitive pose sequence.
@@
Adaptation Strategy. Efficient adaptation strategies of pre-trained models, including adapter-based (Rebuffi et al., 2017), prefix-based (Li & Liang, 2021), and LoRA (Hu et al., 2022) techniques, have been applied to various downstream applications. These strategies achieve advanced performance without fine-tuning all the model parameters. In the field of image synthesis (Rombach et al., 2022), LoRA serves as a plug-and-play extension for the pre-trained model to bring tremendous and specific styles to the images. This work is inspired by this idea and intends to bring the person-specific style to the pre-trained networks that predict facial expressions from speech. A similar idea in speech synthesis (Hsieh et al., 2022) manages to clone the voice of a new speaker with LoRA. Note that we refer to the models proposed in Sec. 3.1 and Sec. 3.2 as the adapters. It is different from the concept of adapter-based strategy Rebuffi et al. (2017) mentioned above.
Remark. Our work differs from previous works by the enhanced data efficiency in style adaptation and overcomes the challenges of catastrophic forgetting for facial expressions and averaged pose generation. To the best of our knowledge, we are the first to simultaneously take facial expressions and head poses into account in speech-driven 3D facial animation.
3 AdaMesh
In this section, we propose AdaMesh, a data-efficient speech-driven 3D facial animation approach tailored for generating personalized facial expressions and head poses. Given the extracted facial expression sequence and head pose sequence from a reference video, together with a driving speech, the expression adapter (Sec. 3.1) is adapted with a specific expression style and outputs rich facial expressions, while the pose adapter (Sec. 3.2) derives a semantic-pose aware style matrix and outputs diverse head poses. The generated facial expressions and head poses are combined by a parametric face model (Li et al., 2017) and rendered to mesh images.
@@ 3 AdaMesh
Figure 1: The overview of expression adapter. (a) The patches on each encoder and decoder denote the MoLoRA parameters which are added to the pre-trained modules after adaptation. (b) A brief illustration of Conformer (Li et al., 2021) (c) demonstrates that MoLoRA is applied to the convolution operator. It is also applied to the linear layers in the proposed expression adapter.
3.1 Expression Adapter
@@
To achieve efficient adaptation for facial expressions, we pre-train the expression adapter to learn general and person-agnostic information that ensures lip synchronization and then optimize the MoLoRA parameters to equip the expression adapter with a specific expression style. As shown in Fig. 1 The expression adapter is composed of an audio encoder, an identity encoder, a style encoder, an expression decoder, and MoLoRA to fine-tune convolutional and linear layers in these modules.
Audio encoder. To obtain robust semantic information from speech, we leverage HuBERT (Hsu et al., 2021), a self-supervised speech representation model, to extract features from raw speech. Unlike Wav2Vec (Baevski et al., 2020) used in previous methods (Fan et al., 2022; Xing et al., 2023), HuBERT can preserve more semantic information and provide better cross-lingual ability. The speech features are downsampled to the same length with expression sequence using convolutional layers with stride 2 and then sent into three Conformer layers (Li et al., 2021).
@@
Identity encoder. Previous methods usually represent the identity with the one-hot label (Xing et al., 2023; Thambiraja et al., 2023) in the animation prediction. They show bad generalization on unseen speaker identity. We utilize the identity parameter extracted by the parametric face model (Li et al., 2017; Feng et al., 2021). The identity sequence in a data sample is averaged along the time dimension to represent the speaker identity and then sent into three convolutional layers with conditional layer normalization (Chen et al., 2021).
@@
Style encoder and Expression Decoder. To provide the talking style which cannot be directly predicted from the semantic speech information, the expression sequence is utilized as a residual condition to the expression prediction. The sequence is compressed to a compact style embedding after two Conformer layers and then concatenated with the encoded speech and identity features into the decoder. The decoder also consists of three Conformer layers, which are stacked in a residual manner rather than cascadation. It is inspired by the residual vector quantization (Zeghidour et al., 2021), which can realize coarse-to-fine prediction.
@@
MoLoRA. We focus on the low-rank training techniques and propose MoLoRA to equip the expression adapter with the inductive bias of new data. As it is known, the vanilla LoRA (Hu et al., 2022) injects trainable rank decomposition matrices into each linear layer of the Transformer architecture. Since the convolutional layer can be viewed as multiple linear transformations with sliding receptive windows, we extend LoRA to the convolution operator for more flexible trainable parameters. The facial expressions contain multi-scale temporal dynamics, where the coarse-grained expression style (e.g., emotion) and fine-grained facial details (e.g., speech-sync lip movements, muscle random wrinkles) inherent in the expression sequence. We empirically found that the vanilla LoRA with only a single rank size filters out the multi-scale dynamics and tends to produce over-smoothed facial expressions. Hence we propose MoLoRA to mix $N$ LoRA with different rank sizes $r_{i}$. Supposing the pretrained weight matrix of 1D convolution $W_{0}\in\mathbb{R}^{m\times n\times k}$, we define the low-rank product $\Delta W^{i}=W_{B}^{i}W_{A}^{i}$, where $W_{B}^{i}\in\mathbb{R}^{m/r_{i}\times k,r_{i}\times k}$ and $W_{A}^{i}\in\mathbb{R}^{r_{i}\times k,n\times r_{i}}$. The weights $W$ after adaptation merge the pre-trained and MoLoRA weights, which can be defined as:
@@
There is a concurrent work (Lialin et al., 2023) that investigates the efficiency in LoRA. However, it stacks multiple LoRA of a single rank size, while MoLoRA focuses on modeling the multi-scale characteristic in facial expressions.
@@
Procedure of the expression adapter. The pretrain, adaptation, and inference procedures of the expression adapter are listed in Alg. 1. Note that the inference stage has a slight difference with the training. The style embedding is obtained from another expression sequence rather than the ground truth. However, this mismatch has no impact on the generation, since the style embedding learns a global talking style for the expression sequence rather than the semantic speech information.
Algorithm 1 Pretrain, Adaptation, and Inference of Expression Adapter.
@@
1:  Pretrain: Train the Expression Adapter $\theta_{E}$ with source training data. The expression sequence for the style encoder is the same with the prediction of the expression adapter.
@@
2:  Adaptation: Update all low-rank matrices $\theta_{MoLoRA}$ for convolution and linear layers in the expression adapter with the adaptation data.
@@
3:  Inference: Merge $\theta_{E}$ and $\theta_{MoLoRA}$. Given a driving speech, and using an expression sequence from the adaptation data as the input for the style encoder, the expression adapter generates expressions that synchronize with the speech and reenact the facial style of the adaptation data.
Figure 2: The overview of pose adapter. (a) Training of VQ-VAE. (b) Pose GPT. (c) The derivation of the semantic-aware pose style matrix and the retrieval strtegy. In the training of the Pose GPT, the pose style embedding is the assigned one-hot label. $S$ denotes the semantic-aware pose style matrix.
3.2 Pose Adapter

### v4

@@ Abstract
Speech-driven 3D facial animation aims at generating facial movements that are synchronized with the driving speech, which has been widely explored recently. Existing works mostly neglect the person-specific talking style in generation, including facial expression and head pose styles. Several works intend to capture the personalities by fine-tuning modules. However, limited training data leads to the lack of vividness. In this work, we propose AdaMesh, a novel adaptive speech-driven facial animation approach, which learns the personalized talking style from a reference video of about 10 seconds and generates vivid facial expressions and head poses. Specifically, we propose mixture-of-low-rank adaptation (MoLoRA) to fine-tune the expression adapter, which efficiently captures the facial expression style. For the personalized pose style, we propose a pose adapter by building a discrete pose prior and retrieving the appropriate style embedding with a semantic-aware pose style matrix without fine-tuning. Extensive experimental results show that our approach outperforms state-of-the-art methods, preserves the talking style in the reference video, and generates vivid facial animation. The supplementary video and code will be available on the project page https://github.com/thuhcsi/AdaMesh.
Index Terms:
Speech-driven animation, talking face generation.
I Introduction
@@
Speech-driven 3D facial animation has been widely explored and attracted increasing interest from both academics and industry. This technology has great potential in virtual reality, film production, and game creation. Most previous works focus on improving synchronization between speech and lip movements [1, 2, 3, 4, 5], however, they have neglected the personalized talking style, including the facial expressions and head poses. The facial animation with expressive facial expressions and diverse head poses contributes to a more vivid virtual character and grants users with a better interaction experience.
Figure 1: The overview of AdaMesh. The expression adapter and pose adapter generate personalized facial expressions and head poses with the given speech signal and reference talking styles.
@@
The recent works [6, 7] attempt to capture the speaking style with a style encoder, but the learned facial expression styles are in lack of sufficient richness. Some works [8, 9, 10] model the person-specific talking style by fine-tuning or adapting modules. In the real application, only a limited amount of video clips, even shorter than 1 minute, are provided by target users to capture the personalized talking style. These methods thus meet several challenges. 1) The few adaptation data probably induces catastrophic forgetting [11] for the pre-trained model and easily leads to the overfitting problem [12]. Specifically for the facial expressions, after the adaptation, the lip synchronization and richness of expressions significantly decrease for the unseen speech inputs. 2) Speech is a weak control signal for the head poses [13, 14]. Adaption with few data or learning mappings on such weak signals leads to averaged generation. The predicted head poses are in lack of diversity and have a smaller variation of movements. 3) The facial expressions and head poses are usually modeled by a one-size-fits-all approach [14, 8], neglecting the unique intrinsic characteristics of facial expressions and head poses.
In this work, to address these challenges, we propose AdaMesh as illustrated in Fig. 1, an adaptive speech-driven 3D facial animation approach. In consideration of the distinct data characteristics of facial expressions and head poses, we devise two specialized adapters—an expression adapter and a pose adapter—tailored to effectively capture the person-specific talking styles, eschewing the one-size-fits-all approach.
@@
The reference facial expression data provided by users are usually scarce. In an effort to capture expression style, directly fine-tuning the entire model with such limited data often leads to severe overfitting or catastrophic forgetting issues. To tackle these problems, we devise an expression adapter built upon the low-rank adaptation technique, LoRA [15], for its high data efficiency in the few-shot model adaptation. To model the multi-scale temporal dynamics in facial expressions, we extend the vanilla LoRA to the convolution operation and further propose a novel mixture-of-LoRA (MoLoRA) strategy, featured by employing LoRA with different rank sizes.
@@
As for the averaged pose generation problem, we recognize it as a repercussion of generating poses via motion regression [4] and a mismatch of the adapted pose style. Hence we formulate the pose generation task as a generative one, for which we employ a vector quantized-variational autoencoder (VQ-VAE) [16] and a generative pre-trained Transformer (PoseGPT) network [17] as the backbone of the pose adapter. The head poses are mostly associated with the conveyance of negation, affirmation, and turnaround in semantics. To balance the strong and weak semantic associations in head pose patterns, we propose the retrieval strategy for the pose style adaptation with a novel semantic-aware pose style matrix.
The main contributions can be summarized as follows:
@@
  • We propose AdaMesh, which separately models the unique data characteristics of facial expressions and head poses and generates personalized speech-driven 3D facial animation with limited adaptation data.
@@
  • We are the first to introduce LoRA to the 3D facial animation task and propose MoLoRA to efficiently capture the multi-scale temporal dynamics in facial expressions.
@@
  • We propose a pose adapter, featured by a semantic-aware pose style matrix and a simple retrieval strategy for diverse pose generation without fine-tuning any parameters.
@@
II Related Work
II-A Speech-Driven 3D Facial Animation
@@
The recent speech-driven 3D facial animation methods can be divided into two categories according to the format of animation. The first category of methods [1, 2, 3, 4, 9] directly predicts vertices of the face from speech and renders the vertices to meshes for visualization. The vertice data is collected through professional 3D scans of human heads [18, 1, 19], which is expensive and time-consuming. The methods trained on such data usually have authentic lip movements. Another category of methods predicts parameters of expressions [20] or blendshapes [21, 22]. These parameters can be reconstructed from the scanned vertices or 2D face videos through a parametric face model [19, 22]. However, the decreasing quality caused by the reconstruction usually leads to lower speech-lip synchronization. The latest methods that consider personalized style mostly adopt expression parameters for style modeling, since the scanned vertices are in the neutral style and lack rich expressions. EmoTalk [22] and EMOTE [20] are related works, which generate emotion-controllable facial animations. EmoTalk adopts unavailable reconstruction pipelines. EMOTE involves a complicated content-emotion disentanglement mechanism. These methods can switch emotions in the training labels but cannot generalize to unseen personalized styles. Mimic [6] and DiffPoseTalk [7] are concurrent works that adopt a style encoder to encode speaking styles, but the generated facial expressions lack sufficient expressiveness. The most related work is Imitator [9], which generates meshes and leverages the expressions from 2D video for style adaptation.
II-B Head Poses Generation
@@
Existing speech-driven 3D facial animation methods mostly neglect the head poses, since the scanned vertices have no head movements. Some portrait-realistic talking face generation methods attempt to model the head poses. A typical way is to copy the head pose sequence from a real video recording [23, 8, 24], which can definitely generate realistic head poses. However, these methods disregard the correlation between speech and head pose, and the repetitive and non-engaging head poses create unnaturalness during long periods of interaction. The rapidly growing research topic, gesture generation from speech [25], also justifies the importance of predicting motions from speech rather than utilizing a template motion sequence. Another way to predict head poses from speech also attracts many works [26, 27, 28, 14, 29, 30, 31]. They have demonstrated the correlation between speech and head pose. However, these methods can only generate head poses with specific styles, and cannot realize flexible style adaptation. Our approach could be regarded as bridging the advantages of the aforementioned two ways. The quantized pose space learned by VQ-VAE ensures realism, while PoseGPT predicts natural but not repetitive poses.
II-C Adaptation Strategy
@@
Efficient adaptation strategies of pre-trained models, including adapter-based [32], prefix-based [33], and LoRA [15] techniques, have been applied to various downstream applications. These strategies achieve advanced performance without fine-tuning all the model parameters. In the field of image synthesis [34], LoRA serves as a plug-and-play extension for the pre-trained model to bring tremendous and specific styles to the images. This work is inspired by this idea and intends to bring the person-specific style to the pre-trained networks that predict facial expressions from speech. Note that we refer to the models proposed in Sec. III-B and Sec. III-C as the adapters. It is different from the concept of adapter-based strategy [32] mentioned above.
II-D Remark
Our work differs from previous works by the enhanced data efficiency in style adaptation and overcomes the challenges of catastrophic forgetting for facial expressions and averaged pose generation. Moreover, compared with the method [10] that learns specific talking styles on the collected dataset, our work intends to transfer unseen personal talking styles and relies on no large-scale expressive facial animation datasets.
@@ II-D Remark
Figure 2: The overview of expression adapter. (a) The patches on each encoder and decoder denote the Conformer blocks are used and MoLoRA parameters are added to the pre-trained modules after adaptation. (b) A brief illustration of Conformer [35] (c) Illustration of 1D-Convolution weights and MoLoRA weights applied on the input features. MoLoRA combines $N$ LoRAs with different rank sizes of $r_{i}$. MoLoRA parameters are added to the convolution and linear layers in the Conformer blocks of the encoders and decoder to efficiently learn the expression style from the reference data.
III AdaMesh
III-A Overview
In this section, we propose AdaMesh, a data-efficient speech-driven 3D facial animation approach tailored for generating personalized facial expressions and head poses. Facial expressions and head poses have distinct and unique data characteristics. The facial expression are strongly correlated with speech, and the basic facial movements can be accurately predicted from the speech content. The expression style can be viewed as a deviation signal superimposed on these movements, resulting in the display of different emotions. The head poses exhibit strong personal styles and convey specific semantics [26, 28]. It can be regarded as a periodic signal superimposed with speech-related peaks and troughs. Therefore, we employ two independent adapters to separately capture expression and head pose styles. Given the facial expression sequence and head pose sequence extracted from a reference video, together with a driving speech, the expression adapter (Sec. III-B) is adapted with the specific expression style and outputs rich facial expressions; while the pose adapter (Sec. III-C) derives a semantic-pose aware style matrix, and outputs diverse head poses with VQ-VAE and PoseGPT. The generated facial expressions and head poses are combined by a parametric face model [19] and rendered to mesh images.
Facial Parameter Representation. We adopt the morphable head model FLAME [19, 36] as 3D face representation. FLAME is described by a combination function $M(\beta,\delta,\psi)$, that takes coefficients describing identity $\beta\in\mathbb{R}^{100}$, expression $\delta\in\mathbb{R}^{50}$, and pose $\psi\in\mathbb{R}^{3k+3}$ ($k=4$ joints) and returns 5023 vertices of xyz dimensions. The 3D vertices can be further rasterized as 2D mesh. More specifically for this work, $\delta$ and the last 3 dimensions of $\psi$ are for facial motion, and the other dimensions of $\psi$ for head pose.
III-B Expression Adapter
@@
Directly fine-tuning the entire model for expression prediction may lead to forgetting previously learned knowledge, i.e., how to generate speech-synced lip movements in this task. To achieve efficient adaptation for facial expressions, we pre-train the expression adapter to learn general and person-agnostic information that ensures lip synchronization and then optimize the MoLoRA parameters to equip the expression adapter with a specific expression style. As shown in Fig. 2 The expression adapter is composed of an audio encoder, an identity encoder, a style encoder, an expression decoder, and MoLoRA to fine-tune convolutional and linear layers in the decoder and the encoders except the audio encoder.
Figure 3: (a) Architecture of the auto-encoder for expression reconstruction. (b) Reconstructed meshes using different rank sizes. (c) Diversity-expression scores for different rank sizes. ①②③ denote rank size 16, 64 and 128.
@@
Audio Encoder & Identity Encoder. To obtain robust semantic information from speech, the self-supervised speech representation model HuBERT [37] is leveraged to extract features from raw speech. These speech features are then sent into three Conformer layers [35]. Previous methods usually represent the identity with the one-hot label [4, 9], leading to the bad generalization on unseen speaker identity. We utilize the identity parameter extracted by the parametric face model [19, 38]. The identity sequence in a data sample is averaged along the time dimension to represent the speaker identity and then sent into three convolutional layers with conditional layer normalization [39].
@@
Style Encoder and Expression Decoder. To provide the talking style that cannot be directly predicted from the semantic speech information, the expression sequence is utilized as a residual condition to the expression prediction. The sequence is compressed to a compact style embedding after two Conformer layers and then concatenated with the encoded speech and identity features into the decoder. The decoder also consists of three Conformer layers, which are stacked in a residual manner [40] rather than cascadation to realize coarse-to-fine prediction.
Figure 4: The overview of pose adapter. (a) Training of VQ-VAE. (b) PoseGPT. (c) The derivation of the semantic-aware pose style matrix and the retrieval strategy. In the training of the PoseGPT, the pose style embedding is the assigned one-hot label for each sample. $S$ denotes the semantic-aware pose style matrix. More details about the training and inference can be found in the supplementary materials.
@@
MoLoRA. We focus on the low-rank training techniques and propose MoLoRA to equip the expression adapter with the inductive bias of new data. As it is known, the vanilla LoRA [15] injects trainable rank decomposition matrices into each linear layer of the Transformer architecture. Since the convolutional layer can be viewed as multiple linear transformations with sliding receptive windows, we extend LoRA to the convolution layer for more flexible trainable parameters. The facial expressions contain multi-scale temporal dynamics, where the coarse-grained global features (e.g., emotion) and fine-grained facial details (e.g., speech-sync lip movements, muscle wrinkles) inherent in the expression sequence. As shown in Fig. 3, we reconstruct the expression coefficients with a simple auto-encoder [41] and inject the vanilla LoRA modules with different rank sizes for expression observation. We empirically found that the LoRA with a small rank size tends to preserve coarse features and a bigger rank size captures more fine-grained dynamics, especially around the cheek region. The quantitative metric in Fig. 3(c) also proves that a bigger rank size results in higher diversity. The latest works [42, 43] demonstrate that the LoRA strategy can efficiently capture facial motion details while some studies [44, 15] show that combining LoRA of different ranks can achieve more efficient training and better performance in large language models. The results in Fig. 3 are in line with the observations of these works.
Hence we propose MoLoRA by mixing $N$ LoRAs with different rank sizes $r_{i}$ for expression style modeling. Supposing the pretrained weight matrix of 1D convolution $W_{0}\in\mathbb{R}^{m\times n\times k}$, we define the low-rank product $\Delta W^{i}=W_{B}^{i}W_{A}^{i}$, where $W_{B}^{i}\in\mathbb{R}^{m/r_{i}\times k,r_{i}\times k}$ and $W_{A}^{i}\in\mathbb{R}^{r_{i}\times k,n\times r_{i}}$. The weights $W$ after adaptation merge the pre-trained and MoLoRA weights, which can be defined as:
@@
We perform the same reshape operation as the vanilla LoRA [15], converting the dimension $\Delta W^{i}$ as $m/r_{i}\times k\times n\times r_{i}\rightarrow{m\times n\times k}$, which is consistent with the pre-trained weights $W_{0}$. There is a concurrent work [44] that investigates the efficiency of LoRA. However, it stacks multiple LoRA of a single rank size, while MoLoRA focuses on modeling the multi-scale characteristic in facial expressions.
@@
Procedure of Expression Adapter. The pipeline of the expression adapter is listed in Alg. 1. Note that the inference stage has a slight difference with the training. The style embedding is obtained from another expression sequence rather than the ground truth. However, this mismatch has no impact on the generation, since the style embedding learns a global talking style for the expression sequence rather than the semantic speech information.
Algorithm 1 Pipeline of Expression Adapter.
@@
  Pretrain: Train the Expression Adapter $\theta_{E}$ with source training data. The expression sequence for the style encoder is the same with the prediction of the expression adapter.
@@
  Adaptation: Update all low-rank matrices $\theta_{MoLoRA}$ for convolution and linear layers in the expression adapter with the adaptation data.
@@
  Inference: Merge $\theta_{E}$ and $\theta_{MoLoRA}$. Given a driving speech, and an expression sequence from the adaptation data as the input for the style encoder, the expression adapter generates expressions that synchronize with the speech and reenact the facial style of the adaptation data.
Loss Function. We adopt the simple mean square error (MSE) loss at both the pretrain and adaptation stages for the expression adapter. The model is optimized for 300k steps and 30 steps at each stage.
III-C Pose Adapter